# Diffusion Autoencoders for Few-shot Image Generation in Hyperbolic Space

### v1

@@ 3.2 Network Architecture
Diffusion Autoencoders. To address the limitations of GAN-based methods, we adopt the approach of DiffAE [55], leveraging a pre-trained Stable Diffusion (SD) model. The SD decoder, $\epsilon_{\theta}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right)$, predicts noise based on timestep $t$ and a high-level semantic condition $\boldsymbol{c}$, which is learned by a semantic encoder mapping input image $x$ to a meaningful latent code $\boldsymbol{c}$. SD also functions as a stochastic encoder that captures a low-level “stochastic” subcode $\boldsymbol{z}_{T}$, as illustrated in Fig. 2.
@@
For high-level semantic capture, we replace the CLIP text encoder with a CLIP image encoder, applying data augmentations (flip, rotation, random crop) to encourage identity-invariant feature learning. To promote diversity, we use only the CLIP image encoder’s class token (dimension $1\times 1024$) for a compact representation, aligning it with the CLIP text feature space via transformer blocks that inject features into the diffusion process through cross-attention. Thus, given reference image $x$, the condition is defined as $\boldsymbol{c}=\operatorname{CLIP}_{\mathrm{I}}\left(x\right)$.
@@
SD is chosen as our diffusion autoencoder for its ability to generate high-quality images and its compatible feature representation with pre-trained CLIP [56]. In this setup, noise is progressively added to $x$ to produce a noisy latent $z_{t}$, where $t$ denotes the number of noise additions. The UNet $\epsilon_{\theta}$ learns to predict the added noise with:
@@
Besides decoding, the SD model can also be used to encode an input image latent $\boldsymbol{z}_{0}$ to the stochastic subcode $\boldsymbol{z}_{T}$ by running its deterministic generative process backward:
$\boldsymbol{z}_{t+1}=\sqrt{\alpha_{t+1}}\mathbf{f}_{\theta}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right)+\sqrt{1-\alpha_{t+1}}\epsilon_{\theta}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right).$ (4)
This process acts as a stochastic encoder, where $\boldsymbol{z}_{T}$ retains only details not captured by $\boldsymbol{c}$, which has limited capacity for stochastic elements. By combining both semantic and stochastic encoders, our diffusion autoencoder captures full image detail and provides a high-level representation $\boldsymbol{c}$ for downstream tasks. Note that the stochastic encoder is omitted during training (Eq. 3) and is used post-training to compute $\mathbf{z}_{T}$ to control the diversity of generated images.
[…]
To learn a semantic hierarchy of real images in hyperbolic space, we minimize the distance between latent codes of images with similar categories and attributes, while distancing codes of dissimilar images. For multi-class classification on the Poincaré disk (Sec. 3.1), we extend multinomial logistic regression (MLR) [20] to hyperbolic space, training an additional linear layer and calculating the hyperbolic loss using negative log-likelihood (NLL) on the softmax output for each class:
@@
where $N$ is the batch size and $p_{n}$ is the probability predicted by the model for the correct class.
To guarantee the network to project back to the CLIP image space, we use L-2 distance reconstruction loss and cosine similarity loss to reconstructed $\boldsymbol{c}^{\prime}$ as same as the original $\boldsymbol{c}$ in CLIP image space:
@@
where $\lambda$ is a trade-off adaptive parameter. This curated set of loss functions ensures the model learns the hierarchical representation and reconstructs images. More details can be found in the supplementary material.
During the inference stage, few-shot image generation can be performed by sampling random latent codes around the latent codes of given reference images that belong to the same “parent” images.
Figure 4: Images with hierarchical semantic similarity generated by HypDAE by sampling “child” images of their “parent” images on FFHQ, where $r_{\mathbb{D}}$ represents the hyperbolic distance from the “parent” images to the center of the Poincaré disk. As the reference image moves from the edge to the center (from fine-grained/certain to abstract/ambiguous), the “children” of the given image become more diverse and less similar to the reference image.
@@
Hyperbolic Delta Mapper. To define editing directions for few-shot image generation, one might use CLIP image features as pseudo-text conditions for text-guided editing in our hierarchical latent space. However, due to the modality gap between image and text feature spaces [44], this approach often results in poorly disentangled manipulations. Inspired by [47], which aligns CLIP text-image features in the $\mathcal{W^{+}}$-space of StyleGAN2 [36], we introduce a Hyperbolic Delta Mapper that aligns CLIP text-image features within hyperbolic space, enabling interpretable latent translation and effective text-driven editing.
In the Hyperbolic Delta Mapper framework, we extract delta image-text features, $\Delta i=i_{2}-i_{1}$, and $\Delta t=t_{2}-t_{1}$, as condition vectors during training and inference. As shown in Fig. 2, training involves obtaining CLIP image space direction $\Delta i=i_{2}-i_{1}$ and corresponding hyperbolic direction $\Delta\boldsymbol{c}_{h}=\boldsymbol{c}_{h2}-\boldsymbol{c}_{h1}$. A latent mapper network in hyperbolic space, the Hyperbolic Delta Mapper, then predicts the editing direction based on:
$\Delta\boldsymbol{c}_{h}^{\prime}=\textit{HypDeltaMapper}\left(\boldsymbol{c}_{h1},i_{1},\Delta i\right),$ (8)
where $\boldsymbol{c}_{h1}$ and $i_{1}$ are used as the input of Hyperbolic Delta Mapper to provide specialized information for the source image.
To train the proposed Delta Mapper, we use the same reconstruction loss as Eq. 6:
$\mathcal{L}_{\text{rec}}(\Delta\boldsymbol{c}_{h},\Delta\boldsymbol{c}_{h}^{\prime})=\|\Delta\boldsymbol{c}_{h}-\Delta\boldsymbol{c}_{h}^{\prime}\|_{2}+1-\cos\left(\Delta\boldsymbol{c}_{h},\Delta\boldsymbol{c}_{h}^{\prime}\right).$ (9)
During inference, text-guided image editing is achieved with the trained Hyperbolic Delta Mapper by replacing $\Delta i$ with $\Delta t=t_{2}-t_{1}$, enabling disentangled feature editing without altering the image’s category.
@@ 4.1 Implementation Details
We choose Stable Diffusion V2.1 [58] as the base generator for the base generative model. During training, we set the image resolution to $512\times 512$. The dimension of the latent code in hyperbolic space is chosen to be $512$. More details can be found in the supplementary material (SM).
@@ 4.2 Datasets
We evaluate our method on Animal Faces [45], Flowers [51], VGGFaces [53], FFHQ [35], and NABirds [67] following the settings described in [16, 41].
@@
DAWSON [43] $3$-shot 188.96 0.0583 208.68 0.0642 137.82 0.0769 181.97 0.1105
F2GAN [31] $3$-shot 120.48 0.2172 117.74 0.1831 109.16 0.2125 126.15 0.2015
WaveGAN [71] $3$-shot 42.17 0.3868 30.35 0.5076 4.96 0.3255 - -
F2DGAN [76] $3$-shot 38.26 0.4325 25.24 0.5463 4.25 0.3521 - -
DeltaGAN [32] $1$-shot 109.78 0.3912 89.81 0.4418 80.12 0.3146 96.79 0.5069
SAGE [16] $1$-shot 43.52 0.4392 27.43 0.5448 34.97 0.3232 19.45 0.5880
HAE [41] $1$-shot 50.10 0.4739 26.33 0.5636 35.93 0.5636 21.85 0.6034
LSO [75] $1$-shot 35.87 0.4338 27.20 0.5382 4.15 0.3834 - -
HypDAE (Ours) $1$-shot 23.96 0.7595 14.31 0.7415 6.25 0.5685 7.64 0.7959
Table 1: FID($\downarrow$) and LPIPS($\uparrow$) of images generated by different methods for unseen categories on four datasets. Bold indicates the best results and underline indicates the second best results. VGGFaces is marked with * because different methods report different numbers of unseen categories on this dataset (e.g. 552 in LoFGAN, 96 in DeltaGAN, 497 in L2GAN, and 572 in HypDAE, HAE, and SAGE).
@@
Figure 9: UMAP visualization of Hyperbolic and Euclidean 2-D embeddings of Animal Faces dataset. We observe that images with clear identities are clustered and positioned near the boundary, while ambiguous samples are located near the center in hyperbolic space. Zoom in to see the details.
@@ 4.3 Analysis of Hierarchical Feature Editing
We examine the properties of the learned hierarchical representations, focusing on how attribute levels correspond to latent code locations in hyperbolic space. As noted in Sec. 3.1, there is a continuum from fine-grained to abstract attributes, corresponding to points from the periphery to the center of the Poincaré disk. We define the hyperbolic radius $r_{\mathbb{D}}$ [^2] The radius of the Poincaré disk in our experiment is about $6.2126$ as the distance of a latent code from the disk’s center. To investigate the effect of $r_{\mathbb{D}}$, we conduct experiments with various radii.
@@
Hierarchical Image Sampling and Interpolation. Fig. 9 shows 2-D embeddings of Animal Faces in the Poincaré disk, visualized via UMAP [48]. By varying the radii of “parent” images, HypDAE controls the semantic diversity of generated images (Fig. 4). Results indicate that identity-relevant attributes change below $r_{\mathbb{D}}\approx 2.0$, while identity-irrelevant attributes vary above $r_{\mathbb{D}}\approx 5.0$, demonstrating that $r_{\mathbb{D}}$ correlates with attribute levels. As $r_{\mathbb{D}}$ decreases, identities become more abstract, resulting in greater diversity among “children” images, as seen in Fig. 11. Smooth interpolation between attributes is achievable in hyperbolic space (Fig. 5) without distortion, as shown in Fig. 6, confirming that HypDAE enables geodesic, hierarchical editing.
@@
Hierarchical Feature Fusion. HypDAE also supports infinite semantic levels for attribute fusion. In Fig. 7, attributes from two images are combined at varying levels along the radius from the disk’s edge to the center, achieving smooth fusion without the limitations of GAN-based generators, which typically support only finite fusion levels (e.g., 18 levels in StyleGAN2’s $\mathcal{W^{+}}$-space).
@@
Ablation Study. Two ablation studies assess the impact of the stochastic encoder and hyperbolic embedding radius on few-shot image quality and diversity. As shown in Fig. 10, the similarity between generated images and a reference image is adjustable via the encoding strength of the stochastic encoder; encoding more information increases similarity but reduces diversity. Additionally, Tab. 2 illustrates how $r_{\mathbb{D}}$ influences image quality and diversity: smaller radii produce higher diversity but can shift category (high-level attributes). Optimal fidelity and diversity balance for few-shot generation is achieved at $r_{\mathbb{D}}\approx 5.5$.
@@
Figure 10: Ablation study on the influence of the encoding strength of the stochastic encoder on four datasets (strength equals $1$ means $x_{0}$ is fully deconstructed, i.e., $x_{T}$ is a Gaussian noise).
@@
Figure 12: Disentangled feature editing with text instructions on four datasets. Zoom in to see the details.

### v2

@@ 3.2 Hyperbolic Diffusion AutoEncoders (HypDAE)
To address the limitations of GAN-based methods, we adopt the approach of DiffAE [47], leveraging a pre-trained SD model. The SD decoder, $\epsilon_{\theta}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right)$, predicts noise based on timestep $t$ and a high-level semantic condition $\boldsymbol{c}$, which is learned by a semantic encoder mapping input image $x$ to a meaningful latent code $\boldsymbol{c}$. SD also functions as a stochastic encoder that captures a low-level “stochastic” subcode $\boldsymbol{z}_{T}$, as illustrated in Fig.˜3.
@@
For the original SD models, the condition $\boldsymbol{c}$ is the given text and is usually processed by a pre-trained CLIP [48] text encoder, outputting 77 tokens. Hence, a naive solution is to directly replace it with CLIP image embeddings to capture high-level semantics. However, this naive solution makes the model easy to remember instead of understanding the context information and copying the content, arriving at a trivial solution. We apply two tricks to avoid this: 1) Strong data augmentations $\mathcal{A}$ (e.g., flip, rotation, blur, and elastic transform) on the reference image $x$ to break down the connection with the source image and encourage identity-invariant feature learning; 2) Content Bottleneck. To promote diversity, we use only the CLIP image encoder’s class token. It compresses the reference image from spatial size $224\times 224\times 3$ to a vector of dimension $1\times 1024$ for a compact representation, aligning it with the original CLIP text feature space of SD via additional fully connected layers (MLP) that inject features into the diffusion process through cross-attention. This mechanism forces the semantic encoder to focus on the main object while ignoring backgrounds and identity-irrelevant features. Thus, given reference image $x$, the condition is defined as: $\boldsymbol{c}=\operatorname{MLP}\left(\operatorname{CLIP}\left(\mathcal{A}\left(x\right)\right)\right)$.
@@
In this setup, noise is progressively added to $\boldsymbol{z}_{0}$ ($x$ encoded by VAE) to produce a noisy latent $\boldsymbol{z}_{t}$, where $t$ denotes the number of noise additions. To utilize the strong image prior learned by SD and CLIP, only the MLP is trainable during the training process. The UNet $\epsilon_{\theta}$ learns to predict the added noise with the condition $\boldsymbol{c}$:
@@
Besides decoding, the SD model can also be used to encode an input image latent $\boldsymbol{z}_{0}$ to the stochastic subcode $\boldsymbol{z}_{T}$ by running DDIM sampling process backward [53]:
$\boldsymbol{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\boldsymbol{z}_{t}+\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)\cdot\epsilon_{\theta}\left(\boldsymbol{z}_{t},t,\boldsymbol{c}\right).$ (4)
For the definition of $\alpha_{t}$ and additional details, please refer to Sec. B of the supplementary material (SM).
This process functions as a stochastic encoder, where $\boldsymbol{z}_{T}$ retains details not captured by the limited-capacity semantic representation $\boldsymbol{c}$. By integrating both semantic and stochastic encoders, the diffusion autoencoder preserves full image details while providing a high-level representation $\boldsymbol{c}$ for downstream tasks. Notably, the stochastic encoder is omitted during training (Eq.˜3) and applied post-training to compute $\boldsymbol{z}_{T}$, enabling control over image diversity.
Hyperbolic Image Encoder. Despite the category-invariant feature extraction in Stage I, the generated image variations remain limited. Thus, Stage II aims to capture hierarchical relationships within the image, enabling latent code editing to enhance diversity while preserving object identity.
To manipulate latent code in hyperbolic space, we need to define a bijective map from $\mathbb{R}^{n}$ to $\mathbb{D}_{c}^{n}$ to map Euclidean vectors to the hyperbolic space and vice versa. A manifold is a differentiable topological space that locally resembles the Euclidean space $\mathbb{R}^{n}$ [34, 35]. For $x\in\mathbb{D}^{n}$, one can define the tangent space $T_{x}\mathbb{D}^{n}_{c}$ of $\mathbb{D}^{n}_{c}$ at $x$ as the first-order linear approximation of $\mathbb{D}^{n}_{c}$ around $x$. Therefore, this bijective map can be performed by exponential and logarithmic maps. Specifically, the exponential map $\exp_{\mathbf{x}}^{c}:T_{\mathbf{x}}\mathbb{D}_{c}^{n}\cong\mathbb{R}^{n}\to\mathbb{D}_{c}^{n}$, maps from the tangent spaces into the manifold. While the logarithmic map $\log_{\mathbf{x}}^{c}:\mathbb{D}_{c}^{n}\to T_{\mathbf{x}}\mathbb{D}_{c}^{n}\cong\mathbb{R}^{n}$ is the reverse map of the exponential map.
We train a hyperbolic image encoder to map latent codes from Euclidean to hyperbolic space using exponential and logarithmic maps at the origin $\mathbf{0}$. Given a latent vector $\boldsymbol{c}\in\mathbb{R}^{1\times 1024}$ from Stage I in CLIP-space, a 5-layer single-head Transformer encoder $\operatorname{E}$ reduces its dimensionality to $\mathbb{R}^{1\times 512}$ before mapping it to hyperbolic space via the exponential map. A hyperbolic feed-forward layer [16] then produces the hierarchical representation $\boldsymbol{c}_{h}$, formulated as: $\boldsymbol{c}_{h}=f^{\otimes_{c}}(\exp_{\mathbf{0}}^{c}(\operatorname{E}(\boldsymbol{c}))),$ where $f^{\otimes_{c}}$ denotes the Möbius linear layer mapping Euclidean to hyperbolic space. To project $\boldsymbol{c}_{h}$ back to the CLIP image space and reconstruct the latent code $\boldsymbol{c}^{\prime}$, we apply a logarithmic map followed by a 30-layer single-head Transformer decoder $\operatorname{D}$: $\boldsymbol{c}^{\prime}=\operatorname{D}(\log_{\mathbf{0}}^{c}(\boldsymbol{c}_{h})),$ which is then fed into the cross-attention layer of the SD model to reconstruct the image $x^{\prime}$.
Loss functions. To learn a semantic hierarchy in hyperbolic space, we minimize the distance between latent codes of similar images while increasing the distance between dissimilar ones. For multi-class classification on the Poincaré disk (Sec.˜3.1), we extend multinomial logistic regression (MLR) [16] to hyperbolic space by incorporating a linear layer and computing the Hyperbolic Loss via the negative log-likelihood (NLL):
@@
where $N$ is the batch size and $p_{n}$ represents the predicted probability of the correct class.
As discussed in Sec.˜3.1, distances in the Poincaré disk grow exponentially with radius. To minimize Eq.˜5, fine-grained image embeddings are pushed toward the disk’s boundary to maximize inter-category distances, while abstract image embeddings—representing shared features across categories—remain near the center.
To ensure that the network projects back to the CLIP image space, we use L-2 distance reconstruction loss and cosine similarity loss to reconstruct $\boldsymbol{c}^{\prime}$ as the same as the original $\boldsymbol{c}$ in CLIP image space:
@@
where $\lambda$ is a trade-off adaptive parameter. This curated set of loss functions ensures the model learns the hierarchical representation and reconstructs images. More details can be found in Sec. A of the supplementary material.
3.3 Pseudo-Labeling
As discussed in Sec.˜1, a key limitation of previous few-shot image generation methods is the reliance on labeled data, which is also evident in our earlier approach (i.e., Eq.˜5). To mitigate this, we leverage a pre-trained CLIP model to predict the pseudo-label. Given an image and a predefined set of class names (e.g., Golden Retriever, Spaniel), we utilize CLIP to compute semantic similarities between the image embedding and class embeddings. The pseudo-label is determined by selecting the class with the highest similarity score, which represents the closest semantic match to the image. This process enables label-free classification by leveraging CLIP’s cross-modal understanding capabilities. See the Sec. A in supplementary material for further details.
3.4 Hyperbolic Latent Editing
To generate diverse images with varying identity-irrelevant features, we edit hyperbolic latent codes via interpolation between two images or by applying random perturbations. In hyperbolic space, the shortest path between two points is defined by the geodesic under the induced distance (Eq.˜2). The geodesic between two embeddings, $\boldsymbol{c}_{hi}$ and $\boldsymbol{c}_{hj}$, is denoted as $\gamma_{\boldsymbol{c}_{hi}\rightarrow\boldsymbol{c}_{hj}}(t)$. For perturbation-based generation, we first rescale the embedding $\boldsymbol{c}_{h}$ of a given image $x_{i}$ to a fixed radius $r_{\mathbb{D}}$, then sample a random vector $\boldsymbol{c}_{hj}$ from seen category embeddings, also constrained to $r_{\mathbb{D}}$. The geodesic between them defines the perturbation direction, enabling diverse image generation. More details and formulas are provided in Sec. C of the appendix.
Figure 5: Images with hierarchical semantic similarity generated by HypDAE by sampling “child” images of their “parent” images, where $r_{\mathbb{D}}$ represents the hyperbolic distance from the “parent” images to the center of the Poincaré disk. As the reference image $x$ moves from the edge to the center (from fine-grained/certain to abstract/ambiguous), the “children” of the reference image become more diverse and less similar to the reference image.
@@ 4.1 Implementation Details
We choose Stable Diffusion V2.1 [50] as the base generator for the base generative model. During training, we set the image resolution to $512\times 512$. The dimension of the latent code in hyperbolic space is chosen to be $512$. More details can be found in the Sec. A of the supplementary material.
@@ 4.2 Datasets
We evaluate our method on Animal Faces [39], Flowers [43], VGGFaces [45], FFHQ [30], and NABirds [58] following the settings described in [14, 37]. Due to the low resolution ($64\times 64$) of the VGGfFces dataset, we use FFHQ [30] to fine-tune the model pre-trained on VGGFaces without supervision and visualize images of human faces with FFHQ.
@@
DAWSON [38] 33-shot 188.96 0.0583 208.68 0.0642 137.82 0.0769 181.97 0.1105
F2GAN [27] 33-shot 120.48 0.2172 117.74 0.1831 109.16 0.2125 126.15 0.2015
WaveGAN [61] 33-shot 42.17 0.3868 30.35 0.5076 4.96 0.3255 - -
F2DGAN [66] 33-shot 38.26 0.4325 25.24 0.5463 4.25 0.3521 - -
DeltaGAN [28] 11-shot 109.78 0.3912 89.81 0.4418 80.12 0.3146 96.79 0.5069
SAGE [14] 11-shot 43.52 0.4392 27.43 0.5448 34.97 0.3232 19.45 0.5880
HAE [37] 11-shot 50.10 0.4739 26.33 0.5636 35.93 0.5636 21.85 0.6034
LSO [65] 11-shot 35.87 0.4338 27.20 0.5382 4.15 0.3834 - -
HypDAE (Euc) 11-shot 25.06 0.7420 20.72 0.7288 6.52 0.5429 8.40 0.7904
HypDAE (Real) 11-shot 23.96 0.7595 14.31 0.7415 6.25 0.5685 7.64 0.7959
HypDAE (Pseudo) 11-shot 24.43 0.7630 13.14 0.7431 5.96 0.5560 7.57 0.7966
Table 1: FID($\downarrow$) and LPIPS($\uparrow$) of images generated by different methods for unseen categories on four datasets. Bold indicates the best results and underline indicates the second best results. VGGFaces is marked with * because different methods report different numbers of unseen categories on this dataset (e.g. 552 in LoFGAN, 96 in DeltaGAN, 497 in L2GAN, and 572 in HypDAE, HAE, and SAGE). We do not compare with diffusion-based models for editing or customization purposes as none of them focus on our task in the same settings.
@@
Figure 10: UMAP visualization of Hyperbolic and Euclidean 2-D embeddings of AnimalFaces dataset. Images with clear identities are clustered and positioned near the boundary, while ambiguous samples are located near the center in hyperbolic space.
@@ 4.3 Analysis of Hierarchical Feature Editing
We examine the properties of the hierarchical representations, focusing on how attribute levels correspond to latent code locations in hyperbolic space. As noted in Sec.˜3.1, there is a continuum from fine-grained to abstract attributes, corresponding to points from the periphery to the center of the Poincaré disk. We define the hyperbolic radius $r_{\mathbb{D}}$ [^2] The radius of the Poincaré disk in our experiment is about $6.2126$ as the distance of a latent code to the disk’s center. To investigate the effect of $r_{\mathbb{D}}$, we conduct experiments with various radii.
@@
Hierarchical Image Sampling and Interpolation. Fig.˜10 shows 2-D embeddings of Animal Faces in the Poincaré disk, visualized via UMAP [40]. By varying the radii of “parent” images, HypDAE controls the semantic diversity of generated images (Fig.˜5). Results indicate that identity-relevant attributes change below $r_{\mathbb{D}}\approx 2.0$, while identity-irrelevant attributes vary above $r_{\mathbb{D}}\approx 5.0$, demonstrating that $r_{\mathbb{D}}$ correlates with attribute levels. As $r_{\mathbb{D}}$ decreases, identities become more abstract, resulting in greater diversity among “children” images, as seen in Fig.˜12. Smooth interpolation between attributes is achievable in hyperbolic space (Fig.˜6) without distortion, as shown in Fig.˜7, confirming that HypDAE enables geodesic, hierarchical editing.
@@
Hierarchical Feature Fusion. HypDAE also supports infinite semantic levels for attribute fusion. In Fig.˜8, attributes from two images are combined at varying levels along the radius from the disk’s edge to the center, achieving smooth fusion without the limitations of GAN-based generators, which typically support only finite fusion levels (e.g., 18 levels in StyleGAN2’s $\mathcal{W^{+}}$-space).
4.4 Ablation Study
@@
Three ablation studies assess the impact of the stochastic encoder and hyperbolic embedding radius on few-shot image quality and diversity. As shown in Fig.˜11, the stochastic encoder regulates the similarity between generated and reference images, where stronger encoding enhances similarity at the cost of diversity. Table˜2 further demonstrates that the hyperbolic embedding radius $r_{\mathbb{D}}$ affects image quality and diversity—smaller radii increase diversity but may alter high-level attributes. The optimal trade-off is observed at $r_{\mathbb{D}}\approx 5.5$. Additionally, Tab.˜3 reveals that identity shifts toward perturbed images when $r_{\mathbb{D}}<4.5$, as measured by CLIP-S (CLIP similarity between generated images and reference images) and CLIP-P (CLIP similarity between generated images and perturbed images).
Euclidean versus Hyperbolic. To evaluate performance gains, we re-trained HypDAE (Euc) in Euclidean space using Eq.˜5. As in Tab.˜1 and Fig.˜13, the hyperbolic space improves performance due to enhanced latent code disentanglement [17]. This is further validated by UMAP visualizations in Fig.˜10. More details in Sec. E of the SM.
Pseudo versus Real Label. The accuracy of the pseudo labels predicted in Sec.˜3.3 for the four datasets are: Animal Faces (48.9%), Flowers (78.6%), VGG Faces (41.5%), and NABirds (39.1%). We re-trained HypDAE (Pseudo) using the pseudo labels. Results in Tab.˜1 indicate that HypDAE (Pseudo) outperforms HypDAE (Real) across most benchmarks, suggesting that noise in human-annotated labels may degrade hierarchical representation learning. This finding underscores that manually labeled data is not essential for few-shot image generation.
@@
Figure 11: Ablation study on the influence of the encoding strength of the stochastic encoder on four datasets (strength equals 11 means $x_{0}$ is fully deconstructed, i.e., $x_{T}$ is a Gaussian noise).
@@
Figure 13: Ablation study of few shot image generation by HypDAE(Hyp) and HypDAE(Euc).